OpenStudy (anonymous):

The number of quadratic equations having real roots and which do not change by squaring their roots? (a)4 (b)3 (c)2 (d)1

OpenStudy (anonymous):

anyone help plz !!!!!!!!!!!!!!

OpenStudy (anonymous):

d

OpenStudy (anonymous):

1^2 = 1^0.5 = 1